Find the slope of the line through the two points (5,3) and (3,2). Enter exact answer as a decimal and show all your work.
Answer by vksarvepalli(154)

slope(m) of a line passing through two points (x1,y1) and (x2,y2) is given by
slope m = (y2-y1)/(x2-x1)
so, slope of the line through the two points (5,3) and (3,2) is
slope = (2-3)/(3-5) = (-1)/(-2)
so slope = 1/2 = 0.5 answer